Given the information below for StartUp.Com, compute the expected share price at the end of 2017 using price ratio analysis. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Year 2013 2014 2015 2016
Price N/A $ 80.12 $ 107.32 $ 116.18
EPS N/A −7.70 −.51 −4.04
CFPS N/A −11.20 −9.40 −4.93
SPS N/A 10.10 13.60 17.10
Answer #1
Calculate Share price
Year 2013 2014 2015 2016
Price 0 80.12 107.32 116.18
EPS 0 -7.70 -51.00 -4.04
CFPS 0 -11.20 -9.40 -4.93
SPS 0 10.10 13.60 17.10
Price/EPS 7.93 7.89 6.79 7.54
SPS Growth 0.35 0.26 0.30
Share price 167.85
Share price 2017 167.85
Calculate Share price
Year 2013 2014 2015 2016
Price 0 80.12 107.32 116.18
EPS 0 -7.7 -51 -4.04
CFPS 0 -11.2 -9.4 -4.93
SPS 0 10.1 13.6 17.1
Price/EPS =C4/C7 =D4/D7 =E4/E7 =(SUM(C9:E9)/3)
SPS Growth =(D7-C7)/C7 =(E7-D7)/D7 =(SUM(D10:E10)/2)
Share price =F9*E7*(1+F10)
Share price 2017 167.85